Consider the following procedure: Pick a number. Add 6 to the number. Multiply the sum by 2. Subtract 12 from the product. Divide the difference by 2. Suppose we pick -8, then what is the resulting number?ย 


1
Expert's answer
2021-11-03T18:14:25-0400

Suppose we pick the number "a." Then the resulting number is


"\dfrac{(a+6)\cdot2-12}{2}=a+6-6=a"

We ontain the same number.

Therefore if we pick -8, then the resulting number is -8.